# Drag Your Gaussian: Effective Drag-Based Editing with Score Distillation for 3D Gaussian Splatting

### v2

@@ 3. Preliminary
3.2. Score Distillation Sampling (SDS).
@@ 4.2. Smooth Geometric Editing
Region-Specific Positional (RSP) Decoder. We introduce the local editing guiding principle in Sec. 4.1, however, the implicit representation of triplane, combined with the latent-based optimization of SDS, inevitably result in changes to regions outside the 3D mask. To address this issue, RSP decoder predict the position shifts $\Delta P$ of the masked 3DGS while introducing a new network to correct the unintended movements in regions outside the 3D masks. Additionally, we propose a regularization loss to further constrain the optimization process. Specifically, The 3DGS $\mathcal{G}$ is divided into two subsets: $\mathcal{G}_{m}=\{({p}_{i},{r}_{i},{s}_{i},o_{i},{c}_{i},{m}_{i})\in\mathcal{G}\mid m_{i}=1\}$ and $\mathcal{G}_{um}=\{({p}_{i},{r}_{i},{s}_{i},o_{i},{c}_{i},{m}_{i})\in\mathcal{G}\mid m_{i}=0\}$. two MLPs, $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, are employed as decoders for the feature $f$ obtained from the MTP encoder. The position shift of each Gaussian primitive $g_{i}$ can be formulated as follows:
@@
Two-stage Dragging. The entire dragging process consists of two stages. The first stage focuses on optimizing the geometric structure, thereby establishing the foundational geometric scaffold of the edited scene. During this stage, the model optimizes the MTP encoder and RSP decoder while freezing all 3DGS parameters and halting the densify and prune operations (Kerbl et al., 2023). As shown in Fig. 10 In the complex deformation task of turning a character’s face, after Stage 1, the Gaussians corresponding to the face are dragged to the target region, while the positions of unmasked Gaussians remain unchanged. The second stage focuses on refining the texture details of the scene. During this stage, other attributes of the 3D Gaussians (color, opacity, rotation, and scale) are primarily optimized, and the densify and prune operations are reactivated. After constructing the basic geometric scaffold in the first stage, the regions outside the desired editing area tend to remain unchanged during the second stage. The newly added 3D Gaussians primarily originate from the splitting or duplication of elements in $\mathcal{G}_{m}$. Consequently, all newly added 3D Gaussians are assigned to $\mathcal{G}_{m}$ for subsequent optimization.
@@ 6.5. Multi-round Dragging
Due to the complexity and diversity of editing tasks, there arises a demand for multi-round dragging. As shown in Fig. 11, we explore the potential of extending DYG to multi-round dragging scenarios. In Fig. 11(a), we perform sequential edits on different targets (e.g., right leg, right arm, left leg), building upon the results of the previous round. Similarly, Fig. 11(b) shows multi-round editing applied to the same target (e.g., gradually raising the man’s right arm). The results illustrate that DYG can be easily adapted to to multi-round dragging scenarios while maintaining notable stability.
@@ A.1. Training Setup
For 3DGS reconstruction, we optimized the Gaussians over 7,000 iterations and set the spherical harmonics to degree 0. In our experiment, batch size is set to 4, and the learning rates for Gaussian’s color, opacity, scale, and rotation are set to $2.5\times 10^{-3}$, $2.5\times 10^{-3}$, $2.5\times 10^{-4}$, and $2.5\times 10^{-3}$, respectively. The shifts in Gaussian position are entirely obtained through MTP encoder and RSP decoder, and we set the learning rate for the MTP encoder was set to $1\times 10^{-3}$, while the learning rate for RSP was $5\times 10^{-4}$. In the first stage, we freeze the Gaussian attributes and only train MTP and RSP. In the second stage, we lower the learning rate of MTP to $1\times 10^{-4}$ to stabilize the scene’s geometric structure and begin training the other Gaussian attributes. For Drag-SDS, we set $\lambda_{\text{lat}}=1$, $\lambda_{\text{img}}=0.1$, and $\lambda_{\text{lora}}=1$, the learning rate of the learnable embedding $\hat{y}_{\emptyset}$ set to $1\times 10^{-3}$, and the lora rank set to 16 with the learning rate of $5\times 10^{-4}$.
A.2. 2 Stage Training
@@ A.4. Total Loss
Figure 13. The MLP structure employed in RSP Decoder.

### v4

@@ 3. Preliminary
3.2. Score Distillation Sampling (SDS)
@@ 4.2. Smooth Geometric Editing
Region-Specific Positional (RSP) Decoder. We introduce the local editing guiding principle in Sec. 4.1, however, the implicit representation of triplane, combined with the latent-based optimization of SDS, inevitably result in changes to regions outside the 3D mask. To address this issue, the RSP decoder predict the position shifts $\Delta P$ of the masked 3DGS while introducing a new network to correct the unintended movements in regions outside the 3D masks. Additionally, we propose a regularization loss to further constrain the optimization process. Specifically, The 3DGS $\mathcal{G}$ is divided into two subsets: $\mathcal{G}_{m}=\{({p}_{i},{r}_{i},{s}_{i},o_{i},{c}_{i},{m}_{i})\in\mathcal{G}\mid m_{i}=1\}$ and $\mathcal{G}_{um}=\{({p}_{i},{r}_{i},{s}_{i},o_{i},{c}_{i},{m}_{i})\in\mathcal{G}\mid m_{i}=0\}$. two MLPs, $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, are employed as decoders for the feature $f$ obtained from the MTP encoder. The position shift of each Gaussian primitive $g_{i}$ can be formulated as follows:
@@
Two-stage Dragging. The entire dragging process consists of two stages. The first stage focuses on optimizing the geometric structure, thereby establishing the foundational geometric scaffold of the edited scene. During this stage, the model optimizes both the MTP encoder and RSP decoder while freezing all 3DGS parameters and halting the densify and prune operations (Kerbl et al., 2023). As shown in Fig. 10 In the complex deformation task of turning a character’s face, after Stage 1, the Gaussians corresponding to the face are dragged to the target region, while the positions of unmasked Gaussians remain unchanged. The second stage focuses on refining the texture details of the scene. During this stage, other attributes of the 3D Gaussians (color, opacity, rotation, and scale) are primarily optimized, and the densify and prune operations are reactivated. After constructing the basic geometric scaffold in the first stage, the regions outside the desired editing area tend to remain unchanged during the second stage. The newly added 3D Gaussians primarily originate from the splitting or duplication of elements in $\mathcal{G}_{m}$. Consequently, all newly added 3D Gaussians are assigned to $\mathcal{G}_{m}$ for subsequent optimization.
@@ 6.5. Multi-round Dragging
Due to the complexity and diversity of editing tasks, there arises a demand for multi-round dragging. As shown in Fig. 11, we explore the potential of extending DYG to multi-round dragging scenarios. In Fig. 11(a), we perform sequential edits on different targets (e.g., right leg, right arm, left leg), building upon the results of the previous round. Similarly, Fig. 11(b) shows multi-round editing applied to the same target (e.g., gradually raising the man’s right arm). The results illustrate that DYG can be easily adapted to multi-round dragging scenarios while maintaining notable stability.
@@ A.1. Training Setup
For 3DGS reconstruction, we optimized the Gaussians over 7,000 iterations and set the spherical harmonics to degree 0. In our experiment, batch size is set to 4, and the learning rates for Gaussian’s color, opacity, scale, and rotation are set to $2.5\times 10^{-3}$, $2.5\times 10^{-3}$, $2.5\times 10^{-4}$, and $2.5\times 10^{-3}$, respectively. The shifts in Gaussian position are entirely obtained through the MTP encoder and the RSP decoder, and we set the learning rate for the MTP encoder was set to $1\times 10^{-3}$, while the learning rate for RSP was $5\times 10^{-4}$. In the first stage, we freeze the Gaussian attributes and only train MTP and RSP. In the second stage, we lower the learning rate of MTP to $1\times 10^{-4}$ to stabilize the scene’s geometric structure and begin training the other Gaussian attributes. For Drag-SDS, we set $\lambda_{\text{lat}}=1$, $\lambda_{\text{img}}=0.1$, and $\lambda_{\text{lora}}=1$, the learning rate of the learnable embedding $\hat{y}_{\emptyset}$ set to $1\times 10^{-3}$, and the lora rank set to 16 with the learning rate of $5\times 10^{-4}$.
A.2. Two-Stage Dragging
@@ A.4. Total Loss
Figure 13. The MLP structure employed in the RSP Decoder.